# DriveX: Driving View Synthesis on Free-form Trajectories with Generative Prior

### v2

@@ Abstract
Driving view synthesis along free-form trajectories is essential for realistic driving simulations, enabling closed-loop evaluation of end-to-end driving policies. Existing methods excel at view interpolation along recorded paths but struggle to generalize to novel trajectories due to limited viewpoints in driving videos. To tackle this challenge, we propose DriveX, a novel free-form driving view synthesis framework, that progressively distills generative prior into the 3D Gaussian model during its optimization. Within this framework, we utilize a video diffusion model to refine the degraded novel trajectory renderings from the in-training Gaussian model, while the restored videos in turn serve as additional supervision for optimizing the 3D Gaussian. Concretely, we craft an inpainting-based video restoration task, which can disentangle the identification of degraded regions from the generative capability of the diffusion model and remove the need of simulating specific degraded pattern in the training of the diffusion model. To further enhance the consistency and fidelity of generated contents, the pseudo ground truth is progressively updated with gradually improved novel trajectory rendering, allowing both components to co-adapt and reinforce each other while minimizing the disruption on the optimization. By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time—unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
@@ Appendix C Additional ablations
Appendix D Comparison with methods incorporating generative prior
Since our work is directly built upon the reconstruction-based driving simulation methods and primarily aims to address the view extrapolation challenge within them, the comparisons in the main paper are mainly against the reconstruction-based approaches. While we acknowledge some concurrent works [44, 20, 28, 4, 41] are sharing a similar scope with ours, most of them have not released their code, making comprehensive and rigorous comparison with them infeasible. Despite that, we try our best to compare with those methods that allow for a reasonable comparison and report the results in Tab. 2. The results indicate that our method performs better than other alternatives, which is also confirmed by the qualitative results in Fig. 4. Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints. Compared to generation-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than that only relying on sparse LiDAR signal. Moreover, our method has significantly faster synthesis speed (34 FPS vs. 0.43 FPS).
Appendix E Application on generated videos
@@ Results
Appendix F Results on the PandaSet dataset
@@
Appendix G More quanlitative results
@@
Figure 10: Failure cases. cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.

### v3

# Driving View Synthesis on Free-form Trajectories with Generative Prior

@@ Abstract
Driving view synthesis along free-form trajectories is essential for realistic driving simulations, enabling closed-loop evaluation of end-to-end driving policies. Existing methods excel at view interpolation along recorded paths but struggle to generalize to novel trajectories due to limited viewpoints in driving videos. To tackle this challenge, we propose DriveX, a novel free-form driving view synthesis framework, that progressively distills generative prior into the 3D Gaussian model during its optimization. Within this framework, we utilize a video diffusion model to refine the degraded novel trajectory renderings from the in-training Gaussian model, while the restored videos in turn serve as additional supervision for optimizing the 3D Gaussian. Concretely, we craft an inpainting-based video restoration task, which can disentangle the identification of degraded regions from the generative capability of the diffusion model and remove the need of simulating specific degraded pattern in the training of the diffusion model. To further enhance the consistency and fidelity of generated contents, the pseudo ground truth is progressively updated with gradually improved novel trajectory rendering, allowing both components to co-adapt and reinforce each other while minimizing the disruption on the optimization. By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time–unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
@@ Appendix C Additional ablations
Appendix D Application on generated videos
@@ Results
Appendix E Results on the PandaSet dataset
@@
Appendix F More quanlitative results
@@
Figure 10: Failure cases. Our method cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.